Of the $50,000 Natasha pocketed on her real estate deal, $20,000 went to charity. She invested part of the remainder in Leaders fund with an annual yeild of 16% and the rest in Templeton Fund with an annual yield of 25%. If she made $6060 on these investments in one year, then how much did she invest in each fund?

Please, kindly help me. I have tried this many times and i could not get the concept to solve the problem. Thank you kindly.

Answer by scott8148(5884)

she made 50k and gave 20k to charity, leaving 30k for investing ... let x=Leaders so 30k-x=Templeton

.16(x)+.25(30000-x)=6060 ... 16x+25(30000-x)=606000 ... 16x+750000-25x=606000 ...-9x=-144000 ... x=16000 ... 30k-x=14k